Two positive charges q1 and q2 are located at the points with radius vectors r1 and r2. Find a negative charge q3 and a position vector r3 of the point at which it has to be placed for the force acting ob each of the three charges to be equal to zero.

Grade:12

1 Answers

Arun
5 years ago
must be 
F1=F12+F13=0 
F2=F12+F23=0 
F3=F13+F23=0 
then 
q2/(R2-R1)^2=q3/(R3-R1)^2 
q1/(R2-R3)^2=q3/(R3-R1)^2 
q1/(R3-R1)^2=-q2/(R3-R2)^2=q2/(R2-R3)^2 
being q1=q2=q 
on find 
(R3-R1)^2=(R2-R3)^2 
from which 
R3=(R1^2-R2^2)/(2R1-2R2) 
and 
q3=q(R3-R1)^2/(R2-R1)^2